# RB-SCD: A New Benchmark for Semantic Change Detection of Roads and Bridges in Traffic Scenes

### v2

@@ III-B Label Imbalance Analysis and Challenges
As shown in Fig. 3, our RB-SCD dataset exhibits a certain degree of sample imbalance In terms of semantic changes, there is an imbalance in the types of changes in the dataset, especially for fine-grained CD tasks. For some types of changes (e.g., road widening, bridge construction), there are more samples, while for other types of changes (e.g., road repair, demolition), there are fewer samples. This imbalance may lead to the model overfitting to the more common change types, while neglecting the rarer ones. Additionally, the imbalance between change and no-change samples is another issue, particularly for CD tasks. The no-change samples are relatively more abundant, which may lead to a decrease in the model’s sensitivity when processing change samples. To improve the accuracy and robustness of the model, we will further collect more samples of different types of changes in future work and employ techniques such as data augmentation and sample weighting to alleviate the model’s difficulty in detecting rare changes.
IV Methodology
IV-A Overview
As shown in Fig. 6, our MFDCD mainly includes two components: (1) DFC module effectively couples high-level and low-level visual features with high-frequency and low-frequency spectral features. (2) TFF module enhances the text information obtained through zero-shot inference, explores the representation of text features in the frequency domain, and promotes the interaction and fusion of multimodal data. The details are as follows.
IV-B DFC Module
Given the bi-temporal images $T_{1}$ and $T_{2}$, ResNet50 [45] is used as the backbone network to extract the corresponding multi-scale features $F_{i}^{t}(i\in\left\{1,2,3,4\right\},t\in\left\{1,2)\right\}$. The features from the first two layers of the network are considered as low-level features, while the features from the last two layers are considered as high-level features. The low-frequency and high-frequency components of these features are then obtained through wavelet transform, which allows for the separation of coarse and fine-grained information in both spatial and frequency domains.
Adaptive Sparse Frequency Fusion: Low-level features are rich in detailed information, while high-frequency features contain information about edges, details, and abrupt changes, which naturally correspond to low-level features. However, high-frequency features are sensitive to noise and may amplify the influence of irrelevant regions. As shown in the upper-right corner of Fig. 6, the ASFF block applies a novel sparsity constraint to the high-frequency features, which is dynamically adjusted during training iterations. By combining the sparsity-constrained high-frequency features with low-level features, both the image’s detailed information (such as edges and textures) and local structural information are preserved. This combination makes the network more responsive to CD, especially in areas with detailed changes. The process is as follows:
$LL_{i}^{t},HL_{i}^{t},LH_{i}^{t},HH_{i}^{t}=\mathcal{W}(F_{i}^{t}),$ (1)
where $t\in\left\{1,2\right\}$, $i\in\left\{1,2,3,4\right\}$, $\mathcal{W}(\cdot)$ represents the discrete wavelet transform. $LL_{i}^{t}$ is the low-frequency band, $HL_{i}^{t}$, $LH_{i}^{t}$, and $HH_{i}^{t}$ signifying the high-frequency bands.
$H_{i}^{t}=Conv(Cat(HL_{i}^{t},LH_{i}^{t},HH_{i}^{t})),$ (2)
$\widetilde{H_{i}^{t}}=\mathcal{S}(H_{i}^{t},\lambda(p),g)=\begin{Bmatrix}H_{i}^{t},&if\left|H_{i}^{t}\right|>g\cdot\lambda(p)\\ 0,&otherwise\end{Bmatrix},$ (3)
$\lambda(p)=\lambda_{0}\cdot e^{-\gamma p},$ (4)
$P_{i}^{t}=Conv(Cat(Up(\widetilde{H_{i}^{t}}),F_{i}^{t})),$ (5)
where $t\in\left\{1,2\right\}$, $i\in\left\{1,2\right\}$. $\mathcal{S(\cdot)}$ denotes the sparsity constraint function, which is used to apply the sparsity operation, $\lambda(p)$ is the dynamic sparsity strength, $\lambda_{0}$ is the initial sparsity strength, $\gamma$ is the sparsity decay rate, $p$ is the current training iteration, and $g$ is a constant controlling the sparsity strength. $H_{i}^{t}$ is the original fused high-frequency feature and $\widetilde{H_{i}^{t}}$ is the high-frequency feature after applying the sparsity constraint. $P_{i}^{t}$ is the enhanced feature that combines high-frequency and low-level information. $Conv(\cdot)$ and $Up(\cdot)$ are convolution and bilinear upsampling operation, respectively.
Bidirectional Temporal Frequency Fusion: As shown in the lower-right corner of Fig. 6. Take the low-frequency features of time $T_{2}$ and the high-level features of time $T_{1}$ as an example, high-level features at time $T_{1}$ typically contain the semantic hierarchy of the image, enabling the expression of global structural changes in the image, such as road widening and bridge construction. Low-frequency features at time $T_{2}$ mainly focus on smooth regions or large-scale change information in the image. By combining high-level features from $T_{1}$ with low-frequency features from $T_{2}$, the network can capture global structural changes across time points while maintaining sensitivity to large-scale, smooth region changes. This models the bidirectional relationship of changes: local changes at $T_{1}$ may lead to semantic-level changes at $T_{2}$, while detailed changes at $T_{2}$ may arise from regional distribution adjustments at $T_{1}$. The details of BTFF block are as follows:
$P_{i}^{1}=Conv(Cat[Up(LL_{i}^{2}),F_{i}^{1}]),$ (6)
$P_{i}^{2}=Conv(Cat[Up(LL_{i}^{1}),F_{i}^{2}]),$ (7)
where $i\in\left\{3,4\right\}$, $P_{i}^{t}$ is the enhanced feature that combines low-frequency information and high-level information.
IV-C TFF Module
@@
As shown in Fig. 6, we leverage the powerful zero-shot inference capability of the CLIP model to obtain textual information from RS images for multimodal modeling. For the categories inferred by the CLIP model, we select the top five categories with the highest probability values. For the foreground features, the text description $T_{d}$ is designed in the format “Remote sensing image at time $\left\{period\right\}$ has a $\left\{value\right\}$ probability of being the $\left\{key\right\}$ ,” where $\left\{period\right\}$ represents either $T_{1}$ or $T_{2}$, $\left\{value\right\}$ is the probability value of the corresponding category and $\left\{key\right\}$ represents the category itself. $E_{T}$ is a pre-trained CLIP text encoder that encodes the text embedding. Considering that discrete Fourier transform can help to reveal high-order semantic patterns in text, especially by identifying more complex semantic information through the descriptors used at different temporal points ($T_{1}$, $T_{2}$). Through Fourier transform, periodic or repetitive structures in the text can be extracted, which helps identify commonalities and differences in changes. We first use Fourier transform to convert the text embeddings from the text domain to the frequency domain:
$\mathbf{T}_{freq}=\mathcal{F}(E_{T}(T_{d})),$ (8)
where $\mathcal{F}(\cdot)$ is the discrete Fourier transform, $\mathbf{T}_{freq}\in\mathbb{C}^{B\times N\times C}$, $B$ represents the batch size, $N$ represents the number of frequency components, $C$ represents the dimension of each frequency component. We introduce a graph filter bank $K=[k_{1},k_{2},...k_{k}]$ for the textual features, which propagate information between the nodes of the graph to filter and refine the frequency information, thereby obtaining salient features related to the changes. For the frequency feature $\mathbf{T}_{freq}$, we split it into real and imaginary parts and concatenate them to construct node features. The feature matrix $\mathbf{X}\in\mathbb{R}^{B\times N\times 2C}$ is the collection of features for all nodes, which allows us to convert frequency domain information into node features for the graph. Assuming that each frequency component is connected to every other frequency component (fully connected graph), the edge set $\mathbf{E}\in\mathbb{Z}^{2\times\frac{N(N-1)}{2}}$ of the graph can be constructed by combining each pair of nodes:
$\mathbf{X}=Cat[\mathfrak{R}(\mathbf{T}_{freq}),\mathfrak{I}(\mathbf{T}_{freq})],$ (9)
$\mathbf{E}=\left\{(i,j)\mid 1\leq i\neq j\leq N\right\},$ (10)
$\mathbf{E}^{(k+1)}=\sigma(\hat{\mathbf{A}}\mathbf{E}^{(k)}\mathbf{W}_{k}),$ (11)
$\mathbf{E}^{f}=\sum_{k=1}^{k}\mathbf{E}^{(k)},$ (12)
where $\mathfrak{R}(\mathbf{T}_{freq})$ and $\mathfrak{I}(\mathbf{T}_{freq})$ represent the real and imaginary parts of the frequency domain features, respectively. $\mathbf{E}^{(k)}$ represents the node features after the $k$-th filter, $\hat{\mathbf{A}}$ is the normalized adjacency matrix, which represents the relationships between nodes, $\mathbf{W}_{k}$ is the weight matrix of the graph convolution filter at the $k$-th layer, and $\sigma$ is the activation function, $\mathbf{E}^{f}$ is the weighted sum of the outputs from multiple filters. We use the text-image encoder $E_{VT}$ of the CLIP model to fuse $\mathbf{E}^{f}$ and the image features. After subtracting the results from the bi-temporal phases, the output is fed into the decoder [46] to obtain our CD results.
IV-D Frequency Transform Function Analysis
Why use discrete wavelet transform for image features and discrete Fourier transform for textual features? The advantage of wavelet transform lies in its ability to analyze images at different scales and locations, making it particularly suitable for multi-level, multi-scale image analysis tasks. As shown in Fig. 2, in remote sensing image CD, wavelet transform helps to extract high-frequency features (e.g., detailed changes) and low-frequency features (e.g., overall structure), aiding in the capture of changes at different spatiotemporal scales in the image. Considering that text is typically discrete sequential data, its changes are often global or frequency-related. Therefore, on the other hand, the Fourier transform can effectively reveal periodic semantic patterns in the text, which contributes to identify patterns and trends in text changes.
V Experiments and Results
V-A Public CD Datasets

### v3

@@ 3.2 Dataset Limitations
As shown in Fig. 4, RB-SCD inevitably exhibits class imbalance across semantic change types. Common transitions (e.g., road widening, bridge construction) are well represented, while rarer cases (e.g., road repair, demolition) appear less frequently. We have used a weighted cross-entropy loss to mitigate the class imbalance issue in the RB-SCD dataset, which assigns higher weights to the losses of under-represented classes. We will further collect more samples of different types of changes in future work and employ techniques such as data augmentation and sample weighting to alleviate the model’s difficulty in detecting rare changes.
Figure 7: Overview of our proposed Multimodal Frequency-Driven Change Detector (MFDCD). The framework takes bi-temporal images ($T_{1},T_{2}$) as input. The Dynamic Frequency Coupler (DFC) processes visual features from a ResNet50 [35] backbone to robustly model linear and heterogeneous transitions. It employs an ASFF block to precisely delineate change boundaries and a BTFF block to model the overall change evolution, preventing fragmentation. Concurrently, the Textual Frequency Filter (TFF) processes CLIP-inferred textual descriptions to resolve high semantic ambiguity. It transforms text into the frequency domain and uses a graph filter to reason about the most relevant semantic concepts. Finally, the enhanced visual features from the DFC, guided by the semantic features from the TFF, are fed into a decoder to generate the final semantic change map.
4 Methodology
4.1 Overview
As shown in Fig. 7, our MFDCD follows an encoder-decoder architecture. Given the bi-temporal images $T_{1}$ and $T_{2}$, we first employ a ResNet50 [35] backbone to extract multi-scale visual features $\mathbf{F}_{i}^{t}\in\mathbb{R}^{B\times C_{i}\times H_{i}\times W_{i}}(i\in\left\{1,2,3,4\right\},t\in\left\{1,2\right\})$. The core innovation consists of two parallel branches: a Dynamic Frequency Coupler (DFC) for linear and heterogeneous semantic change modeling and a Textual Frequency Filter (TFF) for semantic disambiguation. Finally, the enhanced features from these two branches are fused and passed to the decoder to generate the final semantic change map.
4.2 Dynamic Frequency Coupler (DFC)
To address the challenge of linear and heterogeneous semantic change transitions, we first employ the DWT operation. Unlike standard convolutions, as shown in Fig. 2(a), DWT naturally decomposes features into two components that are crucial for this task: high-frequency components, which are sensitive to the detailed boundaries where heterogeneous land-cover transitions occur, and low-frequency components, which are ideal for capturing the long-range continuity and global structure of linear targets like roads. However, merely separating these frequency bands is insufficient; they must be intelligently integrated back into the feature hierarchy. Therefore, as shown in Fig. 7, DFC incorporates two specialized fusion sub-modules: the Adaptive Sparse Frequency Fusion (ASFF) and the Bidirectional Temporal Frequency Fusion (BTFF) block.
Adaptive Sparse Frequency Fusion (ASFF). ASFF block is designed for precise boundary delineation. It achieves this by fusing the distilled high-frequency details with the rich, fine-grained low-level spatial features from the backbone. Low-level features are rich in detailed information, while high-frequency features contain information about boundaries and abrupt changes, which naturally correspond to low-level features. Let $\mathbf{F}_{i}^{t}\in\mathbb{R}^{B\times C_{i}\times H_{i}\times W_{i}}(i\in\left\{1,2\right\},t\in\left\{1,2\right\})$ be the low-level visual features from layer $i$ at time $t$. We first apply the DWT to obtain the frequency sub-bands:
$\mathbf{LL}_{i}^{t},\mathbf{HL}_{i}^{t},\mathbf{LH}_{i}^{t},\mathbf{HH}_{i}^{t}=\mathcal{W}(\mathbf{F}_{i}^{t}),$ (1)
where $\mathbf{LL}_{i}^{t}$ is the low-frequency band, $\mathbf{HL}_{i}^{t},\mathbf{LH}_{i}^{t},$ and $\mathbf{HH}_{i}^{t}$ signifying the high-frequency bands. $\mathcal{W}(\cdot)$ is the DWT operation. The high-frequency sub-bands are then fused into a unified high-frequency representation:
$\mathbf{H}_{i}^{t}=C_{fuse}(Cat(\mathbf{HL}_{i}^{t},\mathbf{LH}_{i}^{t},\mathbf{HH}_{i}^{t})),$ (2)
where $C_{fuse}(\cdot)$ is a convolutional block and $Cat(\cdot)$ is the concatenation operation. However, high-frequency features are highly susceptible to noise. Especially in the early stages of training, when the model is not yet well-optimized, these high-frequency features often contain significant clutter and irrelevant details. Therefore, we design a principled, sparsity-aware gating mechanism. $S(\cdot)$ is defined as a dynamic sparsity constraint function, which adaptively thresholds the high-frequency features based on the training progress:
$\tilde{\mathbf{H}}_{i}^{t}=S(\mathbf{H}_{i}^{t},\lambda(p),g)=\begin{dcases}\medspace\medspace\mathbf{H}_{i}^{t},&{if}\quad|\mathbf{H}_{i}^{t}|>g\cdot\lambda(p)\\ \medspace\medspace\medspace\mathbf{0},&\quad\quad{otherwise}\end{dcases},$ (3)
where $\lambda(p)=\lambda_{0}e^{-\delta p}$ is the dynamic sparsity strength, $\lambda_{0}$ is the initial sparsity strength, $\delta$ is the sparsity decay rate, $p$ is the current training iteration, and $g$ is a constant controlling the sparsity strength. Early in training, a strong constraint (large $\lambda(p)$) allows only the most salient boundary features to pass, ensuring stable learning. As training progresses, the constraint $\lambda(p)$ decays, gradually permitting finer details to be fused for precise boundary refinement. This coarse-to-fine approach is more robust than a fixed or learnable threshold, adapting to the model’s learning stage. Finally, we treat this refined high-frequency map as a spatially-aware gate to modulate the low-level features. The final fused feature representation $\mathbf{P}_{i}^{t}$ is generated by a fusion block $F_{\text{ASFF}}$:
$\mathbf{P}_{i}^{t}=F_{\text{ASFF}}(\mathbf{F}_{i}^{t},\tilde{\mathbf{H}}_{i}^{t})=C_{out}(Cat(Up(\tilde{\mathbf{H}}_{i}^{t}),\mathbf{F}_{i}^{t})),$ (4)
where $Up(\cdot)$ is a bilinear upsampling operator to match spatial dimensions, and $C_{out}$ is the output convolution layer.
Bidirectional Temporal Frequency Fusion (BTFF). To robustly model the change evolution of linear structures and prevent fragmentation, the model must explicitly compare the structural information across two time points. The BTFF module is designed to achieve this via a cross-temporal structure injection mechanism. The core principle is to directly infuse the low-frequency structural representation from one time point into the high-level feature space of the other, thereby forcing the network to reason about semantic features in the context of the opposing temporal structure.
Let $\mathbf{F}_{i}^{t}\in\mathbb{R}^{B\times C_{i}\times H_{i}\times W_{i}}(i\in\left\{3,4\right\},t\in\left\{1,2\right\})$ be the high-level features and $\mathbf{LL}_{i}^{t}$ be the low-frequency structural components at time $t$ for layer $i$. To compute the temporally-informed feature representation $\mathbf{P}_{i}^{1}$ for time $t=1$, we inject the structural information from time $t=2$. The low-frequency component from the second time point, $\mathbf{LL}_{i}^{2}$, represents a structural prior. To prepare this prior for fusion, it must be projected into a higher-dimensional space compatible with the semantic features. This is achieved by a projection function $\Phi_{L}(\cdot)$, composed of a convolutional block followed by an upsampling operator $Up(\cdot)$:
$\widehat{\mathbf{LL}}_{i}^{2}=\Phi_{L}(\mathbf{LL}_{i}^{2})=Up(C_{proj}(\mathbf{LL}_{i}^{2})),$ (5)
where $C_{proj}(\cdot)$ maps the low-frequency features to an intermediate representation.
Next, the semantic feature space of the first time point is augmented by the structural prior from the second. We define an augmentation operator $\Psi(\cdot,\cdot)$ which concatenates the two representations and applies a non-linear fusion transformation $C_{trans}(\cdot)$ to model their complex interactions:
$\mathbf{P}_{i}^{1}=\Psi(\mathbf{F}_{i}^{1},\widehat{\mathbf{LL}}_{i}^{2})=C_{trans}(Cat(\mathbf{F}_{i}^{1},\widehat{\mathbf{LL}}_{i}^{2}))+\mathbf{F}_{i}^{1},$ (6)
where $\mathbf{P}_{i}^{1}$ represents the learned structural difference and contextual information derived from the cross-temporal comparison. The process is performed symmetrically to compute $\mathbf{P}_{i}^{2}$ by injecting $\mathbf{LL}_{i}^{1}$ into $\mathbf{F}_{i}^{2}$, thus creating a bidirectional flow of structural information:
$\mathbf{P}_{i}^{2}=\Psi(\mathbf{F}_{i}^{2},\widehat{\mathbf{LL}}_{i}^{1})=C_{trans}(Cat(\mathbf{F}_{i}^{2},\widehat{\mathbf{LL}}_{i}^{1}))+\mathbf{F}_{i}^{2}.$ (7)
This mechanism ensures that the model’s understanding of each time point is conditioned on the structural reality of the other, making it highly effective at modeling change regions.
4.3 Textual Frequency Filter (TFF)
While recent vision-language models like ChangeCLIP [20] leverage textual priors, their fusion in the noisy spatial-domain still struggles to resolve deep semantic ambiguities. To overcome this, our Textual Frequency Filter (TFF) operates directly in the frequency domain to achieve a more robust semantic alignment. The core idea is to transform textual embeddings into a global signature that separates stable semantic concepts from high-frequency noise, thus providing clearer guidance to disambiguate challenging visual scenes. Our empirical validation in Fig. 8 confirms this insight. We visualize the magnitude spectra of CLIP-generated text embeddings for representative change and no-change scenarios. In the no-change scene (Fig. 8(b)), the $T_{1}$ and $T_{2}$ spectra are nearly identical, demonstrating the stability of the spectral representation. Conversely, in the change scenes (Fig. 8(a) and  8(c)), the spectra diverge significantly, as highlighted by the large spectral difference (gray area). This demonstrates that semantic change is effectively encoded as a quantifiable signal in the frequency domain. Therefore, we employ the DFT as the foundational tool within our TFF.
Figure 8: Visualization of text embedding spectra in the frequency domain by using the discrete Fourier transform. The blue dashed line represents the spectrum of the $T_{1}$ image description, while the red solid line is the $T_{2}$ description. The gray shaded area highlights the absolute difference between the two spectra.
@@ 4.3 Textual Frequency Filter (TFF)
As shown in Fig. 7, we first leverage the powerful zero-shot inference capability of the CLIP [36] model to obtain textual information from RS images for multimodal modeling. For the categories inferred by the CLIP model, we select the top five categories with the highest probability values. For the foreground features, the text description $T_{d}$ is designed in the format “Remote sensing image at time $\left\{period\right\}$ has a $\left\{value\right\}$ probability of being the $\left\{key\right\}$ ” where $\left\{period\right\}$ represents either $T_{1}$ or $T_{2}$, $\left\{value\right\}$ is the probability value of the corresponding category and $\left\{key\right\}$ represents the category itself. We leverage a pre-trained CLIP text encoder [36], denoted as $E_{T}$, to obtain token-level embeddings for each textual description $T_{d}$.
$\mathbf{E}^{t}=E_{T}(T_{d}).$ (8)
This process yields a feature tensor $\mathbf{E}^{t}\in\mathbb{R}^{B\times N\times D}$, where $B$ is the batch size, $N$ is the sequence length (number of tokens), and $D$ is the feature dimension for each token.
To transform this sequence of token features into the frequency domain, we apply the DFT along the sequence dimension $N$. This operator, denoted as $\mathcal{F}$, analyzes the semantic patterns across the entire sequence. For each sample $b$ in the batch and each feature dimension $d$, the $k$-th component of the resulting frequency spectrum $\mathbf{T}_{{freq}}\in\mathbb{C}^{B\times N\times D}$ is computed as:
$\mathbf{T}_{{freq}}(b,k,d)=\mathcal{F}(\mathbf{E}^{t})(b,k,d)=\sum_{n=0}^{N-1}\mathbf{E}^{t}(b,n,d)\cdot e^{-j2\pi kn/N},$ (9)
where $n$ is the index over the sequence of tokens, $n\in\{0,\dots,N-1\}$, $k$ is the index for the frequency components, $k\in\{0,\dots,N-1\}$, $d$ is the index for the feature dimensions, $d\in\{0,\dots,D-1\}$, $j$ is the imaginary unit. The resulting $\mathbf{T}_{{freq}}$ is a complex-valued tensor that encodes how semantic features are distributed across different frequencies of the text sequence. This allows us to capture global relational patterns and periodic structures, yielding a more robust representation for subsequent reasoning.
However, these frequency components are not independent. Complex semantic concepts are often defined by global and potentially non-local relationships across the spectrum. To capture these dependencies, we model the frequency components as nodes and introduce a graph filter bank to act as a semantic reasoning engine. Specifically, we employ a fully connected graph, a design choice that imposes the weakest structural prior. This is crucial because key semantic concepts (e.g., bare land and road) often have strong, non-local relationships within our structured text descriptions, which would be overlooked by models assuming locality, such as sparse graphs based on token adjacency. This maximally flexible structure allows our graph filter bank to learn arbitrary combinations of frequencies that characterize specific change types, without being prematurely constrained by locality assumptions. For the frequency feature $\mathbf{T}_{freq}$, we split it into real and imaginary parts and concatenate them to construct node features. The feature matrix $\mathbf{X}$ is the collection of features for all nodes, which allows us to convert frequency domain information into node features for the graph:
$\mathbf{X}=Cat(e(\mathbf{T}_{freq}),Im(\mathbf{T}_{freq})),\mathbf{X}\in\mathbb{R}^{B\times N\times 2C},$ (10)
where $Re(\mathbf{T}_{freq})$ and $Im(\mathbf{T}_{freq})$ represent the real and imaginary parts of the frequency domain features, respectively. The edge set $\mathbf{E}$ can be constructed by combining each pair of nodes:
$\mathbf{E}=\left\{(i,j)\mid 1\leq i\neq j\leq N\right\},\mathbf{E}\in\mathbb{Z}^{2\times\frac{N(N-1)}{2}}.$ (11)
The graph filter bank is denoted as $\mathcal{K}=\{k_{1},k_{2},\dots,k_{L}\}$ and implemented as a set of $L$ parallel layers, $\{\mathcal{G}_{l}\}_{l=1}^{L}$, where each layer $\mathcal{G}_{l}(\cdot)$ corresponds to a filter $k_{l}$. Unlike a 1D CNN that captures only local dependencies, a graph filter performs relational reasoning over the entire spectral structure. Its message-passing mechanism is designed to learn a filter function that enhances or suppresses global spectral patterns associated with specific semantic changes. Each filter then is parameterized by its own weight matrix $\mathbf{W}_{l}$ and learns to capture a specific type of semantic relationship from the node features $\mathbf{X}$, thereby obtaining salient features related to the changes:
$\mathbf{X}^{\prime}_{l}=\mathcal{G}_{l}(\mathbf{X},\hat{\mathbf{A}})=\sigma(\hat{\mathbf{A}}\mathbf{X}\mathbf{W}_{l}),$ (12)
where $\mathbf{X}^{\prime}_{l}$ represents the node features after the $l$-th filter, $\hat{\mathbf{A}}$ is the normalized adjacency matrix derived from $\mathbf{E}$ and $\sigma$ is the activation function. The outputs from all filters in the bank $\mathcal{K}$ are then aggregated to form a comprehensive, refined semantic representation $\mathbf{X}_{T}$:
[…]
Let $V_{ct}$ be the spatial visual context tokens from the image backbone. A context decoder ${E}_{\text{VT}}$, aligns the refined frequency-domain semantics $\mathbf{X}_{T}$ with this spatial visual evidence to compute a visually-informed feature:
$\mathbf{X}_{T}^{refined}={E}_{\text{VT}}(V_{ct},\mathbf{X}_{T}).$ (14)
By analyzing textual semantics in the frequency domain and aligning them with spatial visual features, TFF effectively resolves semantic ambiguities and provides reliable guidance for change prediction.
5 Experiments and Results
5.1 CD Datasets